A police jeep is chasing a culprit going on a motarbike .the motar bike crosses a turning at a speed of 72 km/h. The jeep follows it as a speed of 90km/h . crossing the turning 10 sec later than the bike. Assuming that they travel at constant speeds, how far from the traning will the jeep catch a with the bike.

Grade:11

1 Answers

Vikas TU
7 years ago
After turning point let the culprit travelled a distance of x meter in 10 seconds.
therefore,
x = 72000/3600 * 10 = 200 meter.
and the jeep distance,
y  = 90000/3600 * 10 = 250 meter.
in order to catch the culprit,
the distnace the jeep would cover =>
let (z+ 200) = 25*t.............(1)
for culprit,
z = 20*t............................(2)
solve the eqns. and get z value.